# Insight Over Sight? Exploring the Vision-Knowledge Conflicts in Multimodal LLMs

### v1

@@ 2.1 Automated Pipeline
Figure 3: Illustration of generating questions across three difficulty levels, from easy to hard based on the conflict target (Action in this example.)
4. Generate Questions.
As illustrated in Figure 3, with the constructed counter-commonsense triplets, the pipeline generates questions across three difficulty levels, from easy to hard:
  • Multiple-Choice Question: For these questions, the templates include “What is/are [Subject] doing [Place]?” and “Where is/are [Subject] [Action]?” for counter-commonsense Action and Place, respectively. For a predetermined option count $m$, our framework selects ($m-1$) seemingly relevant but incorrect targets from the candidates. To this end, the framework partitions candidates into $m-1$ bins $\{B_{1},...,B_{m-1}\}$ based on their NPMI scores relative to the context pair. We randomly sample one candidate from each bin, together with the counter-commonsense target, to formulate the question’s multiple-choice options.
  • Subjective Question: The question text in the multiple-choice question is reused, but no list of options is included. To guide the model toward generating a brief and specific response, we add the constraints like “Answer with a single phrase”, with slight adjustment according to the specific prompt for different MLLMs.
2.2 ConflictVis Benchmark Construction
Table 1: Statistics of the constructed benchmark.
@@ 2.2 ConflictVis Benchmark Construction
We introduce how to use the automated pipeline to construct the ConflictVis benchmark. We utilize top 100K sentences from the Open Mind Common Sense (OMCS) dataset (Singh et al., 2002) as our knowledge base. Developed by the MIT Media Lab, OMCS is a rich collection of human commonsense knowledge, consisting of assertive sentences contributed by thousands of volunteers. From this dataset, we extract the 100 most frequent Subject phrases ($N_{S}=100$), along with 150 most frequent phrases for Actions and Places ($N_{A}=N_{P}=150$). Using feedback from the LLM Vicuna-1.5-13b, [^3] https://huggingface.co/lmsys/vicuna-13b-v1.5 we select the top 3 phrases with the highest NPMI scores to create context pairs for each subject ($K=3$). In addition, we choose the top 3 targets with the lowest NPMI scores to assemble the candidate set of triplets ($M=3$). After manually filtering out unexpected combinations, the remaining triplets are used for image generation. The low-quality images are again manually removed. The specific guidelines and results in the quality control process are elaborated in the Appendix B. Following this two-stage generation and filtering process, the statistics of the resulting benchmark are detailed in Table 1. In total, ConflictVis consists of 374 original images with 1,122 QA pairs spanning two conflict targets and three difficulty levels, all verified by human experts. Figure 10 in Appendix shows the instances for the subject “a musician”.
@@ 3.1 Setup
To explore the behavior of MLLMs when encountering commonsense knowledge conflicts, we perform a comprehensive evaluation on 9 MLLMs with 7 representative open-source MLLMs ranging from 8B to 34B, and 2 state-of-the-art commercial MLLMs. This evaluation covers the following five model series: LLaVA (8B, 13B, 34B) (Liu et al., 2024b; c), BLIP-2 (12.1B, 13B) (Li et al., 2023b; Dai et al., 2023), Qwen-VL (9.6B) (Bai et al., 2023), GPT-4o (OpenAI, 2024b) and Claude-3.5-Sonnet (Anthropic, 2024).
@@
We use Accuracy and Memorization Ratio (MR) (Longpre et al., 2021) as the main evaluation metrics in our experiment. Both metrics require classifying the model’s responses into different categories. For Yes-No and Multiple Choice questions, where there is a unique answer, we use exact matching for classification. For Subjective questions, due to the open-ended nature and the complexity of the task (i.e., evaluating both textual and visual relevance and correctness), we rely on human annotators to perform the classification (Guidelines and details are included in Appendix C).
@@ 3.2 Sanity Test
where the context and target are filled with the specific phrases from the input. For example, the query for the instance in Figure 3 is “Based on common sense, is it possible for the baby on the bed fixing a computer?” The model’s responses are categorized as “negate” (indicating a conflict), “accept” (indicating no conflict), and “other” (where the model does not give a direct answer). The results, shown in Figure 4, demonstrate that the vast majority of cases in ConflictVis indeed present valid knowledge conflicts that can be identified by MLLMs, with conflict rates ranging from 80.5% for GPT-4o to 99.2% for LLaVA-1.5-13B.
Figure 5: Model uncertainty on traditional VQA dataset and on our benchmark.
Is ConflictVis a more challenging benchmark for MLLMs?
Unlike the traditional VQA benchmarks, where visual information typically aligns with the model knowledge, our ConflictVis presents conflicts between these two sources of knowledge, making it a more challenging benchmark for MLLMs. To empirically validate our claim, we conduct a comparative analysis of model uncertainty between our ConflictVis and the conventional VQA benchmark (Antol et al., 2015). A higher degree of uncertainty in model predictions on a particular benchmark indicates that the benchmark poses a greater challenge for the model to navigate. Figure 5 shows the results, where higher entropy value denotes more uncertainty. It is evident that the model’s predictions are more uncertain on ConflictVis than the standard VQA benchmark, highlighting the increased challenge posed by our ConflictVis benchmark.
@@ 3.3 Benchmarking MLLMs
(a) Yes-No Questions
(b) Multi-Choice Questions
(c) Subjective Questions
Figure 6: MLLMs’ response distributions on different question types. (Responses in Vision category are considered correct.)
In this section, we assess how well the MLLMs handle conflicts between vision information and parametric knowledge using our ConflictVis benchmark. Following Longpre et al. (2021), we compare model predictions (i.e., responses) with and without the inclusion of the image and then quantify the instances where model’s predictions are influenced by the provided image:
  • Align with its predictions without the given image (Knowledge, $P_{K}$): This outcome suggests that the model generates its answer based on its pre-existing parametric knowledge rather than the visual information presented.
  • Yield a correct answer that aligns with the visual information (Vision, $P_{V}$): This indicates that the model is capable of adapting to new information, adjusting its output to match the visual input.
  • Result in a different answer altogether (Other): This demonstrates that the model can modify its output based on visual input, though the result does not perfectly align with the visual information.
Figure 6 presents the results. Ideally, the model should output the Vision answer, supported by visual information, rather than the Knowledge answer derived from textual training, or any Other answers. However, both open-source and commercial MLLMs often revert to producing the Knowledge answer, ignoring the visual information, to varying degrees. In general, commercial MLLMs perform better than the open-source counterparts, particularly in handling subjective questions. Among the open-source MLLMs, LLaVA-NeXT-34B, being the largest model, delivers the best overall performance, achieving an average prediction accuracy of 84.6%. The top commercial model, GPT-4o, reaches an average prediction accuracy of 89.9%.
Figure 7: Memorization ratio (MR) of MLLMs, where higher MR values indicate greater model overstability on parametric knowledge.
MLLMs are more likely to overly rely on parametric knowledge for simpler questions.
One interesting observation from Figure 6 reveals that MLLMs exhibit poorer performance on straightforward Yes-No questions. Notably, the open-source BLIP-2-12B achieved a low accuracy of 39.3%, while the commercial Claude-3.5-Sonnet managed only 56.4%.
We leverage Memorization Ratio (MR) as a metric to empirically estimate the model’s overstability — it’s brittleness to changing information (Longpre et al., 2021):
@@
Figure 7 illustrates the MR across different question types. Notably, the MR values for Yes-No questions are significantly higher than for the other two question types across all MLLMs, indicating that MLLMs tend to overly rely on parametric knowledge when answering simpler questions.
(a) Counter-commonsense Action
(b) Counter-commonsense Place
Figure 8: MLLMs’ response distributions on two distinct categories of conflict targets.
@@
Our ConflictVis evaluates two distinct types of conflict targets: counter-commonsense actions and places. As shown in Figure 8, the results show that counter-commonsense actions pose more of a challenge than places do. Specifically, the average prediction accuracy for counter-commonsense actions stands at 73.9%, significantly lower than the 85.2% recorded for places. Similarly, the MR for actions is markedly higher at 23.9%, compared to 13.4% for places. This discrepancy could be attributed to the richer visual context available for identifying places (e.g., “in a studio”) as opposed to the sparse cues for actions (e.g., “paddling the boat”).
3.4 Steering MLLMs Towards Vision
Figure 9: Input-output relevance in a failure case where visual information is underutilized.
Underutilization of visual information.
Our primary experiment demonstrates that when MLLMs encounter visual inputs that contradict their pre-existing knowledge, around 18% of the responses generated are not aligned with the visual context but rather reflect the model’s prior knowledge. To further understand this discrepancy, we analyze the overall input-output relevancy inside MLLMs (Stan et al., 2024), seeking insights that could guide improvement strategies. Figure 9 illustrates a common case in which the model LLaVA-1.5-13B failed to accurately respond to a subjective question, highlighting the challenge. The model pays more attention to the textual query than to the visual context. Accordingly, the attended tokens {“musician”, “doing”, “studio”} guide the model to generate the answer “playing guitar” based on the parametric knowledge.
Table 2: MLLMs’ accuracy under different improvement methods.
Model Y-N MC Subj. Avg.
Qwen-VL-Chat-9.6B 69.8 80.5 89.3 79.9
[…]
     + FoV Prompting (this work) 82.4 83.2 87.4 84.3
@@ 3.4 Steering MLLMs Towards Vision
     + FoV Prompting (this work) 75.9 96.5 98.9 90.5
Effectiveness of Existing Methods.
Based on the above observation, one straightforward improvement strategy is to enhance the impact of visual context when generating answers. We consider several training-free approaches along this direction and propose a novel prompting method:
  • Visual Contrastive Decoding (Leng et al., 2024) contrasts the output distributions derived from original and distorted visuals to ensure the generated content closely adheres to the visual inputs.
  • Pay Attention to Image (Liu et al., 2024f) adaptively adjusts and amplifies the attention weights assigned to image tokens, thereby giving greater prominence to visual elements.
  • Chain-of-Thought (CoT) prompting (Wei et al., 2022b) enables complex reasoning capabilities through intermediate reasoning steps.
While Wei et al. (2022b) demonstrates that CoT can enhance the commonsense reasoning capabilities of LLMs, its applicability to MLLMs in leveraging visual elements under conflict scenarios may not be directly transferable. In response to this problem, we propose a simple and effective prompting technique named Focus-on-Vision (FoV) prompting to explicitly instruct MLLMs to prioritize visual information, which can be seamlessly incorporated into current MLLMs:
[textual query] Please focus on the visual information.
Table 2 shows the results. We adhere to the methodologies outlined in the original papers by implementing Visual Contrastive Decoding (VCD) and Pay Attention to Image (PAI) techniques on the Qwen-VL-Chat-9.6B and LLaVA-1.5-13B models. Both approaches significantly improve the models’ performance, aligning with the outcomes reported in the literature for enhancing MLLMs trustworthiness. Interestingly, we notice that CoT prompting diminishes the performance of the models. Upon examining the failure cases, we find that though CoT envokes the commonsense reasoning abilities of MLLMs, these models can over-rely on their parametric knowledge to draw the conclusion, which results in an increased number of inaccurate response or refusals. We present the corresponding failure cases in Appendix E.2. In contrast, our Focus on Vision (FoV) prompting technique addresses this issue by guiding MLLMs to prioritize visual information, which boosts performances across models.
4 Related Work
Knowledge Conflicts.
Knowledge conflicts in LLMs can be divided into three categories (Xu et al., 2024b): within the retrieved context (Chen et al., 2022), within model’s parametric knowledge (Huang et al., 2023), and between the context and the model’s parametric knowledge (Xie et al., 2024; Wu et al., 2024a). These conflicts can lead to incorrect or inconsistent responses, undermining the model’s trustworthiness (Xu et al., 2024b; Xie et al., 2024). In MLLMs, these conflict types expand to multimodal inputs, where conflicts can occur between the input image and the text instruction (Liu et al., 2024e; Han et al., 2024), or when the image contains counterfactual information that contradicts the model’s parametric knowledge (Liu et al., 2024e; Guan et al., 2024).
To evaluate the conflicts between visual information and the parametric knowledge in MLLMs, HallusionBench (Guan et al., 2024) manually edits collected informational graphics to create normal-counterfactual image pairs, and verifies model consistency with simple Yes-No questions. AutoHallusion (Wu et al., 2024b) introduces an automated approach to generate counterfactual scenarios by altering correlated objects in images, utilizing Yes-No questions to probe object existence and spatial relationships. In the context of commonsense knowledge, PhD (Liu et al., 2024e) generates counter-commonsense images through manual collection and synthesis, employing Subjective questions for assessment. Our proposed ConflictVis distinguishes itself by:
  • Incorporating a wider range of question types and conflict targets, including Yes-No, Multiple-Choice, Subjective questions and counter-commonsense actions and places. This approach allows for a more thorough evaluation of MLLMs. Our findings indicate that MLLMs tend to disregard visual information more frequently when responding to simpler questions and conflicting action is more challenging for MLLMs to resolve.
  • Automatically generating counter-commonsense images based on the co-occurrence of knowledge components and crafting questions around these counter-commonsense aspects, ensuring better image-question alignment. With the aid of advanced text-to-image models, our framework can generate instances tailored to specific needs.
Hallucination in MLLMs. Hallucination in MLLMs refers to cases where the model generates descriptions that conflict with the given visual context. MLLMs’ hallucinations are typically categorized based on the type of incorrect information, such as non-existent objects, incorrect object attributes, and inaccurate object relations (Liu et al., 2024a). Relevant research has mainly focused on two key areas: developing benchmarks and metrics to detect hallucinations (Rohrbach et al., 2018; Li et al., 2023c; Liu et al., 2023a; Wang et al., 2023) and proposing strategies to mitigate them (Liu et al., 2023a; Leng et al., 2024; Huang et al., 2024; Liu et al., 2024f).
In the context of knowledge conflicts, hallucination occurs when the model gives precedence to its intrinsic knowledge rather than the visual context (Liu et al., 2023a; Guan et al., 2024; Liu et al., 2024e; a). Our study explores this issue by simulating knowledge conflicts that MLLMs may face during inference, and evaluates the effectiveness of different strategies to mitigate hallucinations (Leng et al., 2024; Liu et al., 2024f) by enhancing the model’s fidelity to visual inputs. Furthermore, we introduce a novel prompting technique that explicitly instructs MLLMs to prioritize visual information, thereby significantly and consistently reducing the occurrence of hallucinations.
Benchmarks for MLLMs. Traditional vision-language benchmarks are designed to assess independent skills, including image captioning (Chen et al., 2015), visual grounding (Rohrbach et al., 2016), and visual question answering (Antol et al., 2015; Hudson & Manning, 2019). However, with the emergence of MLLMs, there is a growing need for more comprehensive and tailored benchmarks. The strong zero-shot abilities and advanced language generation skills exhibited by MLLMs can render traditional benchmarks insufficient, as they may not account for the diversity of responses or the full range of MLLM capabilities. To address these limitations, researchers have developed more complex and comprehensive benchmarks to assess MLLMs across a wider range of capabilities (Yue et al., 2024; Fu et al., 2023; Li et al., 2023a; Liu et al., 2023c). Meanwhile, diagnostic benchmarks have also been proposed to evaluate specific challenges or traits in MLLMs, such as hallucination (Li et al., 2023c; Liu et al., 2023a), social bias (Howard et al., 2024), and model safety (Liu et al., 2023b). ConflictVis is a pioneering analytical benchmark to introduce conflicting visual contexts against the model’s commonsense knowledge, allowing us to investigate model performance in the face of vision-knowledge conflicts.
@@ 5 Conclusion
This paper introduces an innovative approach to understand and mitigate the significant challenge of vision-knowledge conflicts in MLLMs, with a particular focus on commonsense knowledge. By developing an automated pipeline and the ConflictVis benchmark, we offer a tool for systematically evaluating and understanding the nuances of these conflicts. Our extensive experiments across nine representative MLLMs reveal a tendency toward over-reliance on parametric knowledge, especially in simpler queries, which underscores the importance of the visual context in these models’ decision-making processes. Building on these insights, we evaluate three improvement methods and design a new Focus-on-Vision prompting to reinforce the significance of visual input, thereby reducing the impact of conflicts. Our work paves the way for future research aimed at improving the robustness and reliability of MLLMs. Our goal is to ensure these models can effectively integrate and interpret multimodal data, mitigating internal knowledge inconsistencies.
Appendix A Benchmark Details

### v2

@@ 3.2 ConflictVis Benchmark Construction
Table 2: Statistics of the constructed benchmark.
@@ 4.1 Setup
To explore the behavior of MLLMs when encountering vision-knowledge conflicts, we perform a comprehensive evaluation on 9 MLLMs including 7 representative open-source MLLMs ranging from 8B to 34B, and 2 state-of-the-art commercial MLLMs. This evaluation covers the following five model series: LLaVA (8B, 13B, 34B) (Liu et al., 2024b, c), BLIP-2 (12.1B, 13B) (Li et al., 2023b; Dai et al., 2023), Qwen-VL (9.6B) (Bai et al., 2023), GPT-4o (OpenAI, 2024b) and Claude-3.5-Sonnet (Anthropic, 2024). The diversity of model architectures and parameter sizes helps to enhance the generalizability of experiment results.
@@
We use Accuracy and Memorization Ratio (MR) (Longpre et al., 2021) as the main evaluation metrics in our experiment. Both metrics require classifying the model’s responses into different categories. For Yes-No and Multiple-Choice questions, where there is a unique answer, we use exact matching for classification. For Open-Ended questions, due to the complexity of the task (i.e., evaluating both textual and visual relevance and correctness) and the instability of LLM evaluation (Stureborg et al., 2024), we rely on human annotators to perform the classification.
Figure 3: MLLM response distributions on different question types.
@@ 4.2 Sanity Test
where the context and target are filled with the specific phrases from the input. For example, the query for the instance in Figure 2 is “Based on common sense, is it possible for the waitress in the kitchen signing a bill?” The model’s responses (“Yes”, “No”, and others) are categorized as “accept” (indicating no conflict), “negate” (indicating a conflict), and “other” (where the model does not give a direct answer). The results, shown in Figure 4, demonstrate that the vast majority of cases in ConflictVis indeed present valid knowledge conflicts that can be identified by MLLMs, with conflict rates ranging from 80.5% for GPT-4o to 99.2% for LLaVA-1.5-13B.
Figure 4: MLLM responses to sanity test inputs.
@@ 4.3 Benchmarking MLLMs
In this section, we assess how well the MLLMs handle conflicts between visual information and parametric knowledge using our ConflictVis benchmark. Following Longpre et al. (2021), we compare model predictions (i.e., answers) with and without the inclusion of the image input and quantify the extent to which the model’s predictions are influenced by the presence of the image. Specifically, we compare the predictions in the current experiment, where the image is provided, to those obtained without the image using the sanity test prompts, and classify the current predictions as:
  • Aligned with the prediction without the given image (Knowledge, $P_{K}$): This outcome suggests that the model generates its answer mainly based on its parametric knowledge rather than the visual information presented.
  • Aligned with the visual information (Vision, $P_{V}$): This indicates that the model is capable of adapting to new information, adjusting its output to match the visual input. Such outputs are categorized as the correct answer.
  • A different answer altogether (Other): This demonstrates that the model can modify its output based on visual input, though the result does not perfectly align with the visual information.
Figure 3 presents the results. Ideally, the model should only output the Vision answer, supported by visual information, rather than the Knowledge answer derived from textual training, or any Other answers. However, both open-source and commercial MLLMs revert to producing the Knowledge answer, ignoring the visual information, to varying degrees. In general, commercial MLLMs perform better than the open-source counterparts, particularly in handling open-ended questions. Among the open-source models, LLaVA-NeXT-34B achieves the highest prediction accuracy at 84.6%, while the top commercial model, GPT-4o, reaches 89.9%. Moreover, our analysis reveals that MLLMs’ uncertainty in answer generation is significantly higher on our ConflictVis compared to the classic VQA benchmark, indicating a greater challenge to the model performance (See Appendix C.2).
Figure 5: Memorization ratio of MLLMs on different question types.
MLLMs are more likely to overly rely on parametric knowledge for Yes-No questions.
One interesting observation from Figure 3 reveals that MLLMs exhibit poorer performance on straightforward Yes-No questions. Notably, the open-source BLIP-2-12B achieves a low accuracy of 39.3%, while the commercial Claude-3.5-Sonnet manages only 56.4%.
We leverage Memorization Ratio (MR) (Longpre et al., 2021) as a metric to empirically estimate the the extent to which the model adheres to its parametric knowledge in the presence of conflicting information.
@@
Figure 5 illustrates the MR across different question types. Notably, the MR values for Yes-No questions are significantly higher than the other two question types across all MLLMs, indicating that MLLMs tend to overly rely on parametric knowledge when answering Yes-No questions. We hypothesize that this could be due to the format of Yes-No questions, which directly present counter-commonsense expressions (e.g., “Is the waitress in the kitchen signing a bill?”) to the MLLMs, while other question types have a less direct phrasing. This counter-commonsense query may discourage the model from thoroughly analyzing the visual input, causing it to immediately output a negation.
@@
Our ConflictVis evaluates two distinct types of conflict targets: counter-commonsense actions (e.g., the waitress in the kitchen signing a bill) and places (e.g., a doctor conducting an experiment at the theater). As shown in Figure 6, counter-commonsense action problems pose more of a challenge than places do. Specifically, MLLMs consistently exhibit a lower accuracy on counter-commonsense action problems, with an average output accuracy of 73.9%, notably lower than the 85.2% recorded for place problems. Similarly, the average Memorization Ratio for action problems is markedly higher at 23.8%, compared to 13.4% for place problems. This discrepancy could be attributed to the richer visual context typically available for identifying places (e.g., numerous background elements suggesting the location “at the theater”), in contrast to the relatively sparse and fine-grained visual cues required to recognize actions (e.g., subtle hand posture for “signing a bill”). Detailed results for each model can be found in the Appendix C.4.
Figure 6: MLLM accuracy on Counter-Commonsense (CCS) Action problems and CCS Place problems.
4.4 Mitigating Vision-Knowledge Conflicts
Figure 7: Failure case analysis of input-output relevancy. The visual information is underutilized.
Underutilization of visual information
Our primary experiment demonstrates that when MLLMs encounter visual inputs that contradict their pre-existing knowledge, around 18% of the responses generated are not aligned with the visual context but rather reflect the model’s prior knowledge. To further investigate this discrepancy, we leverage the tool proposed by Stan et al. (2024) to analyze input-output relevancy scores, which identify the most relevant parts of the input to the generated output based on model attention. Figure 7(a) illustrates a case in which the model LLaVA-1.5-13B fails to accurately answer an open-ended question, representing a common failure pattern. In general, the model assigns greater weight to the textual input than to the visual context, resulting in an inaccurate response. In this case, the most attended tokens, “musician”, “doing”, “street”}, guide the model to generate “playing guitar” based on its parametric knowledge, while the image tokens, denoted by the <image> symbol in the bar plot, play a less significant role in answer generation. The image-output relevancy maps in Figure 7(b) further indicate that the model may not effectively leverage visual information during answer generation. In such cases, the model is more susceptible to vision-knowledge conflicts, as the visual input is largely underutilized. (More cases can be found in Appendix D.1.)
Effectiveness of Existing Methods
Based on the above observations, a straightforward strategy for improvement is to enhance the influence of visual context during the answer generation process. We consider several improvement approaches:
  • Visual Contrastive Decoding (VCD) (Leng et al., 2024) contrasts the output distributions derived from original and noisy visuals to ensure the generated content adheres to the visual inputs.
  • Pay Attention to Image (PAI) (Liu et al., 2024f) adaptively adjusts and amplifies the attention weights assigned to image tokens, thereby giving greater prominence to visual elements.
  • Vision-Centric Reasoning (VR) employs structured intermediate reasoning steps to thoroughly analyze visual inputs, achieved by Chain-of-Thought (CoT) prompting (Wei et al., 2022b) or supervised finetuning on the long CoT multimodal reasoning dataset (Xu et al., 2024b).
Furthermore, we propose a direct prompting technique, termed Focus-on-Vision (FoV) prompting, which explicitly instructs MLLMs to prioritize visual information. This can be seamlessly integrated into current MLLMs using the following template: [textual query] Please focus on the visual information.
Model YN MC OE Avg.
@@ 4.4 Mitigating Vision-Knowledge Conflicts
     + FoV 75.9 96.5 98.9 90.5
Table 3: MLLM accuracy under different improvement methods. YN: Yes-No, MC: Multiple-Choice, OE: Open-Ended.
Table 3 presents the results of the evaluated improvement methods. For VCD and PAI, we follow the methodologies outlined in the original papers and apply each technique to the corresponding applicable MLLMs, and observe performance improvements for both methods. To implement Vision-Centric Reasoning (VR), we explore two approaches: (1) applying Chain-of-Thought (CoT) prompting, and (2) fine-tuning models on the multimodal reasoning dataset introduced by Xu et al. (2024b). However, the results show minimal performance gains from finetuning, and vanilla CoT prompting can even degrade model accuracy. Upon analyzing failure cases, we find that teaching MLLMs to reason in a straightforward, step-by-step manner can lead to increased reliance on textual knowledge at test time. For example, our case study (in Appendix D.2) suggests that during CoT reasoning, models tend to rationalize the inputs using their existing commonsense knowledge. However, counter-commonsense inputs are inherently difficult to rationalize or explain. As the output text accumulates, it may become increasingly difficult for the model to accurately interpret the visual input given its own generated explanations. This often results in self-contradictory answers or refusals to answer. In contrast, our Focus-on-Vision (FoV) prompting mitigates this issue by explicitly directing the model’s attention to the visual input and consistently boosts model performance. Nevertheless, none of the approaches can entirely resolve the vision-knowledge conflict, especially for open-source models. For instance, LLaVA-1.5-13B and Qwen-VL-Chat still exhibit notable error rates of 13.2% and 15.7%, respectively, even when equipped with the most effective improvement method. This highlights the profound challenge posed by vision-knowledge conflict.
@@ 5 Conclusion
In this paper, we develop an automated framework and construct the ConflictVis benchmark to systematically evaluating and understanding the nuances of vision-knowledge conflicts in MLLMs at the commonsense level. Our experiments across nine representative MLLMs reveal a tendency toward over-reliance on parametric knowledge, especially among Yes-No and action-related questions. We further evaluate three improvement methods and design a new prompting technique to mitigate the conflicts, but the problem remains notable. Collectively, our insights lay the groundwork for future research to develop more reliable MLLMs.
Acknowledgment
This paper was supported by the Guangdong Basic and Applied Basic Research Foundation (No. 2024A1515010145) and the Shenzhen Science and Technology Program (Shenzhen Key Laboratory Grant No. ZDSYS20230626091302006).
Limitations
In this study, we focused on evaluating the reliably of multimodal large language models (MLLMs) under vision-knowledge conflict. While our work provides insights into the model performance and effectiveness of various mitigation approaches, it’s important to acknowledge certain limitations: (1) Limited root cause analysis: while we conducted failure case analysis using relevancy maps, the fundamental causes of the unexpected behaviors exhibited by MLLMs under vision-knowledge conflicts remain underexplored. We hypothesize that these behaviors may stem from biases in the training data, where MLLMs are predominantly aligned with image-text pairs that conform to commonsense expectations encoded during text pretraining, with limited exposure to counter-commonsense scenarios. To further validate this hypothesis, future work could involve controlled experiments comparing models trained with and without curated datasets designed for counter-commonsense alignment, in order to investigate the resulting behavior differences. (2) Unique probability model: due to budget constraints, our framework relies on Vicuna-1.5-13B to estimate co-occurrence probabilities, which may introduce model-specific features into the benchmark construction. Although our quality control processes help mitigate this issue, future work can explore more balanced probability estimation approaches, such as averaging probabilities across multiple models.
Ethics Statements
While our benchmark is designed to be safe and useful for research purposes, the developed framework for generating counter-commonsense inputs warrants careful consideration regarding its potential applications. In certain contexts, the framework may produce images that deviate from conventional norms, and if misused, could result in cases that some users may perceive as harmful or offensive. To mitigate such risks, we recommend implementing precautionary measures, including the use of a carefully curated and pre-filtered input corpus, as well as incorporating a manual verification process. Additionally, we encourage the use of text-to-image generative models through their official platforms and in accordance with their terms of use, which can help effectively leverage built-in safety mechanisms and reduce the likelihood of generating inappropriate or harmful content.
Appendix A Benchmark Construction Details